OpenStudy (anonymous):

What indicators predict that a quadratic funtion will have a complex solution? [Solve and show all work] x^2 - 10x = -13 x = -b +- squareroot b^2 -4ac OVER 2a

OpenStudy (irishboy123):

focus on "b^2 -4ac " as you need to take its square root.....

OpenStudy (anonymous):

Well, I'm not sure how to solve it. Can you teach me? @IrishBoy123

OpenStudy (irishboy123):

what is b and what is c in your original quadratic?

OpenStudy (irishboy123):

this is the quadratic you posted: \(\large x^2 - 10x = -13 \) and this is the additional formula you also posted: \(\large x = \frac{-b \ \pm \ \sqrt{ b^2 -4ac} }{ 2a}\)

OpenStudy (irishboy123):

maybe it would be better to start by asking you what you are trying to do to the quadratic? i know but do you also know?

OpenStudy (anonymous):

No, I'm unsure on how to solve this question...

OpenStudy (irishboy123):

that's my question....you are trying to find the value of x where the quadratic = 0, is that right?

OpenStudy (anonymous):

Well, the question is asking to find complex solutions... Which I dont really understand what it means by that.

OpenStudy (anonymous):

and I guess it's asking to solve also.. So yeah. solving for x

OpenStudy (irishboy123):

take your original equation \(x^2 - 10x = -13 \) and re-write it as \(x^2 - 10x +13 =0\) , yep?!! then compare it to this \(a x^2 + b x + c = 0\) and then apply that second equation you posted, the one with "\(b^2 - 4 ac\)" in it [i set it out again above] that second equation gives you the roots (x intercepts) of the quadratic, ie x values for when quadratic = 0 try it and see where you go.....

OpenStudy (irishboy123):

clearly if \(b^2 - 4ac \lt 0\) you might have an issue (ie complex roots) but run with it first....

OpenStudy (anonymous):

Well, I'm sorta confused on how to plug it in and solve it... Can you help me on that?

OpenStudy (anonymous):

@IrishBoy123 you there?

OpenStudy (irishboy123):

compare \(x^2−10x+13=0\) to \(ax^2+bx+c=0\) so \(a = 1, b = -10, c = 13\) plug these into \(\huge x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\)

OpenStudy (irishboy123):

what do you get?

OpenStudy (anonymous):